OpenStudy (anonymous):

I need help. Methane (CH4), ammonia (NH3), and oxygen (O2) can react to form hydrogen cyanide (HCN) and water according to this equation: CH4 + NH3 + O2  HCN + H2O You have 8 g of methane and 10 g of ammonia in excess oxygen. How many grams of hydrogen cyanide will be formed? Show your work.

OpenStudy (anonymous):

2CH4 + 2NH3 + 3O2 --> 2HCN + 6H2O

OpenStudy (anonymous):

0 g of NH3 is 0.588 moles (10/17) and 8 g of CH4 is 0.5 moles (8/16).

OpenStudy (anonymous):

You have excess oxygen, so you need to find the limiting reagent

OpenStudy (anonymous):

0.588/2 = 0.294, and 0.5/2 = 0.25. Therefore, CH4 is your limiting reagent, and this is what you need to use to find the amount of HCN produced.

OpenStudy (anonymous):

CH4 and HCN have the same coefficient in the equation

OpenStudy (anonymous):

If you're using up all 0.5 moles of CH4, you are producing 0.5 moles of HCN. 0.5 moles of HCN is 13.5 g (0.5 * 27).

OpenStudy (anonymous):

So the Balanced equation : 2CH4 + 2NH3 +302 --> 2HCN + 6H20

OpenStudy (anonymous):

CH4=8/16=0.5 NH3=10/17=0.59 <- which is the number of moles

OpenStudy (anonymous):

One mole of CH4 requires one mole of NH3.

OpenStudy (anonymous):

There are 0.59 moles of NH3

OpenStudy (anonymous):

0.5*27= 13.5 g.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

yes, thank you very much!

OpenStudy (anonymous):

No problem!